A nurse has two solutions that contain different concentrations of a certain medication. One is a 20% concentration and the other is a 5% concentration. How many cubic centimeters of each should he mix to obtain 10 cc of a 8% solution
Answer by Alan3354(69443)

Do it like this one:
=========================================
A large solar heating panel requires 110 gallons of a fluid that is 40% antifreeze. The fluid comes in either a 65% solution or a 20% solution. How many gallons of each should be used to prepare the 110-gallon solution?
====================
s = amount of 65%
t = amount of 20%
----
s + t = 111 (total solution)
65s + 20t = 40*110 (total antifreeze)